# Autoregressive Moving-average Attention Mechanism for Time Series Forecasting

### v1

@@ Abstract
We propose an Autoregressive (AR) Moving-average (MA) attention structure that can adapt to various linear attention mechanisms, enhancing their ability to capture long-range and local temporal patterns in time series. In this paper, we first demonstrate that, for the time series forecasting (TSF) task, the previously overlooked decoder-only autoregressive Transformer model can achieve results comparable to the best baselines when appropriate tokenization and training methods are applied. Moreover, inspired by the ARMA model from statistics and recent advances in linear attention, we introduce the full ARMA structure into existing autoregressive attention mechanisms. By using an indirect MA weight generation method, we incorporate the MA term while maintaining the time complexity and parameter size of the underlying efficient attention models. We further explore how indirect parameter generation can produce implicit MA weights that align with the modeling requirements for local temporal impacts. Experimental results show that incorporating the ARMA structure consistently improves the performance of various AR attentions on TSF tasks, achieving state-of-the-art results. The code implementation is available at the following link.
@@ 1 Introduction
In recent years, autoregressive (AR) decoder-only Transformer-based models (Vaswani, 2017; Radford, 2018) have been widely used in sequence modeling tasks across fields such as NLP (Brown et al., 2020; Touvron et al., 2023), CV (Chen et al., 2020; Esser et al., 2021; Chang et al., 2022), and audio (Borsos et al., 2023). This structure is well-suited for various sequential generation and prediction tasks. However, in typical sequence modeling tasks like time series forecasting (TSF), there has been less exploration of this architecture compared to other structures. Most of the best-performing recent TSF models are encoder-only Transformers (Liu et al., 2024a; Nie et al., 2022), MLPs (Das et al., 2023; Lu et al., 2024), or even linear models (Zeng et al., 2023; Xu et al., 2024). The few relevant discussions mainly focus on using pretrained autoregressive LLMs or similar structures for few-shot and zero-shot prediction (Gruver et al., 2023; Jin et al., 2024; Das et al., 2024; Liu et al., 2024b), with little research directly evaluating their TSF performance in end-to-end training. Therefore, this paper will first briefly demonstrate that with appropriate tokenization and training methods, a basic AR Transformer is enough to achieve results comparable to the state-of-the-art (SOTA) baselines, as shown in Fig. 1 and Fig. 2.
Recently, efficient linear AR attention variants have been explored and developed (Katharopoulos et al., 2020; Hua et al., 2022), reducing the time complexity of standard softmax attention from $O(N^{2})$ to $O(N)$. Researchers have found that adding a gating decay factor or a similar exponential moving average (EMA) structure to AR structure, as in gated linear attention (Ma et al., 2022; Yang et al., 2024), enhances linear attention’s ability to model local patterns and improves performance. The success of these approaches inspired us to introduce a more comprehensive full autoregressive moving-average (ARMA) structure into existing AR attention mechanisms and explore the performance of the ARMA Transformers in TSF.
@@
We propose the ARMA attention mechanism, which integrates a moving-average (MA) term into various existing AR attention mechanisms. Our method improves the TSF performance of AR Transformers without significantly increasing computational costs, maintaining $O(N)$ time complexity and original parameter size. We design an indirect MA weight generation to obtain the MA output without explicitly computing the MA attention matrix, preserving efficiency of linear attentions. We explore specific techniques for generating implicit MA weights to ensure proper decoupling and handling of short-term effects. Extensive experiments and visualization analyses demonstrate that ARMA balances long- and short-term dependencies, significantly improving AR Transformers and achieving state-of-the-art TSF results.
@@
a) We demonstrate that, with appropriate tokenization and preprocessing methods, an AR Transformer is enough to achieve the level of existing SOTA baselines. Furthermore, the introduction of ARMA attention enables the decoder-only Transformer to outperform SOTA baselines.
b) We propose the ARMA attention mechanism, which introduces an MA term into existing AR attention without increasing time complexity or parameter size. By adding the MA term to various AR attention mechanisms, the resulting ARMA Transformers significantly improve forecasting performance compared to their AR counterparts.
@@ 2.2 Appropriate tokenization for autoregressive forecasting
To prevent error accumulation, we use an autoregressive Transformer (Fig. 1) that treats one-step prediction as the complete forecast. Inspired by PatchTST (Nie et al., 2022), we adopt a channel-independent approach, predicting each series separately and applying RevIN (Kim et al., 2022) to each. For an input series of length $L_{I}$ in $\mathbf{S}_{I}$, we apply non-overlapping patches with a patch size $L_{P}$, dividing the input into $N=\frac{L_{I}+P}{L_{P}}$ patches, where $P$ is zero-padding for divisibility. This ensures that each out-of-sample prediction token covers the entire forecasting length $L_{P}$, thereby avoiding error accumulation.
Figure 2: Box plots of performance rankings from 48 sub-experiments across 12 datasets. Green represents ARMA Transformers, yellow AR Transformers, and red the baselines, with triangles indicating mean rankings. AR Transformers perform comparably to baselines, while ARMA Transformers significantly outperform their AR counterparts. See Table 2 and 8 for more details.
@@ 2.3 Preliminaries: decoder-only Transformer
$\begin{gathered}\text{Attn}(\mathbf{X})=\text{softmax}\left(\mathbf{M}\odot(\mathbf{Q}\mathbf{K}^{\top})\right)\mathbf{V}\mathbf{W}_{o},\ \ \text{with}\ \ \mathbf{Q},\mathbf{K},\mathbf{V}=\mathbf{X}\mathbf{W}_{q},\mathbf{X}\mathbf{W}_{k},\mathbf{X}\mathbf{W}_{v}\\ \mathbf{X}:=\mathbf{X}+\text{Attn}(\text{LN}(\mathbf{X})),\ \ \text{then}\ \ \mathbf{X}:=\mathbf{X}+\text{MLP}(\text{LN}(\mathbf{X}))\end{gathered}$ (1)
@@ 2.4 Preliminaries: efficient linear attention mechanisms
2.5 ARMA attention mechanism
Figure 3: ARMA attention structure with the indirect MA weight generation method applied to softmax and linear attention. See Table 1 for more calculation details.
In these attention mechanisms, the next-step prediction at time $t$ is a weighted sum of all previous values $\bm{v}_{i}\in\mathbb{R}^{1\times d}$, with weights $\mathbf{w}_{t,i}\in\mathbb{R}^{1\times d}$ derived from interactions between $\bm{q}_{t}$ and $\bm{k}_{i}$. Naturally, we can write these attention mechanisms in an AR model structure:
@@ 2.5 ARMA attention mechanism
$\bm{v}_{t+1}=\bm{o}^{\text{AR}}_{t}+\bm{o}^{\text{MA}}_{t}+\bm{\epsilon}_{t}=\sum_{i=1}^{t}\mathbf{w}_{t,i}\odot\bm{v}_{i}+\sum_{j=1}^{t-1}\bm{\theta}_{t-1,j}\odot\bm{\epsilon}_{j}+\bm{\epsilon}_{t},\ \ \bm{r}_{t}=\sum_{j=1}^{t-1}\bm{\theta}_{t-1,j}\odot\bm{\epsilon}_{j}+\bm{\epsilon}_{t}$ (2)
The structure of the MA output $\bm{o}^{\text{MA}}_{t}=\sum_{j=1}^{t-1}\bm{\theta}_{t-1,j}\odot\bm{\epsilon}_{j}$ resembles the AR term and could potentially be computed using an attention mechanism. For simplicity, we consider a single channel of the $d$-dimensional space, with other channels an be handled in parallel. We express the matrix form of the $\bm{r}_{t}$ in Eq. 2 for one channel as:
$\begin{gathered}\begin{pmatrix}r_{1}\\ r_{2}\\ \vdots\\ r_{t}\end{pmatrix}=\begin{pmatrix}0&0&\cdots&0&0\\ \theta_{1,1}&0&\cdots&0&0\\ \theta_{2,1}&\theta_{2,2}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \theta_{t-1,1}&\theta_{t-1,2}&\cdots&\theta_{t-1,t-1}&0\end{pmatrix}\begin{pmatrix}\epsilon_{1}\\ \epsilon_{2}\\ \vdots\\ \epsilon_{t}\end{pmatrix}+\begin{pmatrix}\epsilon_{1}\\ \epsilon_{2}\\ \vdots\\ \epsilon_{t}\end{pmatrix}\\ \mathbf{r}=(\mathbf{I}+\mathbf{\Theta})\epsilon,\ \ \mathbf{\epsilon}=(\mathbf{I}+\mathbf{\Theta})^{-1}\mathbf{r},\end{gathered}$ (3)
where $\mathbf{\Theta}$ is a strictly lower triangular matrix of MA weights for this channel. Once the attention mechanism determines $\bm{o}^{\text{AR}}_{t}$ and $\bm{\theta}_{t-1,j}$, we can calculate $\bm{r}_{j}=\bm{v}_{j+1}-\bm{o}^{\text{AR}}_{j}$ (token shifting) for all $j\leq N-1$ and determine $\epsilon_{j}$ via matrix inversion. Substituting these back into Eq. 2 yields the final ARMA attention output $\bm{o}^{\text{AR}}_{t}+\bm{o}^{\text{MA}}_{t}$ for step $t$.
@@
2.6 Indirect MA weight generation
We need an approach that can leverage linear attention’s efficiency to compute $\bm{o}^{\text{MA}}_{t}$ without the costly $\mathbf{\Theta}^{N\times N}$ matrix operations. Instead of separately calculating attention weights to determine $\bm{\epsilon_{j}}$ as value input and recomputing the whole MA output, we aim to use a linear RNN to collect all keys and values at once. We observe from Eq. 3 that there is already a sequential relationship between $\bm{r}_{j}$ and $\bm{\epsilon}_{j}$, and $\bm{r}_{j}$ can be computed directly once $\bm{o}^{\text{AR}}_{t}$ is determined. Therefore, we implicitly compute the MA weights of $\bm{\epsilon_{j}}$ by using $\bm{r}_{j}$ as value input for the MA component instead of $\bm{\epsilon}_{j}$. Let $\bm{\beta}_{t-1,j}$ denote the generated attention weights corresponding to $\bm{r}_{j}$ at step $t$, and let $\bm{\theta}_{t-1,j}$ here be the implicit MA weights hiddenly linked to the generated $\bm{\beta}_{t-1,j}$. Based on Eq. 3, we establish:
$\begin{gathered}\sum_{j=1}^{t-1}\bm{\beta}_{t-1,j}\odot\bm{r}_{j}=\sum_{j=1}^{t-1}\bm{\theta}_{t-1,j}\odot\bm{\epsilon}_{j}\Leftrightarrow\mathbf{B}\mathbf{r}=\mathbf{\Theta}\mathbf{\epsilon}\ \ \text{(for one channel)}\\ \mathbf{B}=\mathbf{\Theta}\cdot(\mathbf{I}+\mathbf{\Theta})^{-1},\ \ \mathbf{\Theta}=\mathbf{B}\cdot(\mathbf{I}-\mathbf{B})^{-1}\end{gathered}$ (4)
$\theta_{ij}=b(1+b)^{i-j-1},\quad\text{where}\ i>j$ (5)
We tested various activation function combinations for $\phi_{q}^{\text{MA}}(\cdot)$ and $\phi_{k}^{\text{MA}}(\cdot)$ to generate $\bm{\beta}_{t-1,j}=\phi_{q}^{\text{MA}}(\bm{q}^{\text{MA}}_{t-1})\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})$ values, as shown in Fig. 4, 7, and 8. We used the sigmoid function $\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})=\sigma(\alpha\bm{k}^{\text{MA}}_{j}/\sqrt{d})$ to obtain values between 0 and 1, where $\alpha=0.05$  [^1] In the key activation, $\alpha$ controls the variance of each row in the $\mathbf{B}$ matrix, indirectly influencing the amount of long-term information (lower left) in the MA weights $\mathbf{\Theta}$. Increasing $\alpha$ would make the MA weights focus more on modeling long-term information. However, since we want the AR weights to handle the long-term component, we set $\alpha$ to a relatively small value. This explains why the rows of the $\mathbf{B}$ matrix appear smooth in the visualization. Refer to Fig. 7 for more details on $\alpha$, and see Fig. 8 for the effects of reversed positive $\phi_{q}$. and $\sqrt{d}$ are scaling factors to maintain small absolute values. Then, we selected a function $\phi_{q}^{\text{MA}}(\cdot)$ to make the product negative. We ultimately chose $\phi^{\text{MA}}_{q}(\bm{q}^{\text{MA}}_{t})=-\text{LeakyReLU}(-\bm{q}^{\text{MA}}_{t}/\sqrt{d})$ with a negative slope of 0.02. The inner negative sign maintains directional consistency (for later parameter sharing), and the outer negative sign encourages a negative output.
(a) $\phi_{q}^{\text{MA}}$: -LeakyReLU, $\phi_{k}^{\text{MA}}$: $\sigma$
(b) $\phi_{q}^{\text{MA}}$: -ReLU, $\phi_{k}^{\text{MA}}$: $\sigma$
(c) $\phi_{q}^{\text{MA}}$: -$\sigma$, $\phi_{k}^{\text{MA}}$: $\sigma$
(d) $\phi_{q}^{\text{MA}}$: -Swish, $\phi_{k}^{\text{MA}}$: $\sigma$
Figure 4: Visualization of the $\mathbf{B}(\textrm{left})-\mathbf{\Theta}(\textrm{right})$ relationship with different $\phi(\cdot)$. We construct the simulated $\mathbf{B}$ matrices using randomly sampled $\bm{q}$ and $\bm{k}$ ($N=64$, $d=32$) from the normal distribution, and display the corresponding implicit $\mathbf{\Theta}$ matrices.
Fig. 4 shows that LeakyReLU provides a balanced lag weight pattern. Unlike ReLU and Sigmoid, which only output values of the same sign, LeakyReLU offers some relaxation while keeping most values negative. This adds flexibility by enabling the desired negative smoothing effect of the MA term, with occasional positive values to enhance modeling flexibility.
To summarize the ARMA attention process with indirect MA weight generation: First, we compute all $\bm{o}^{\text{AR}}_{t}$ using the selected attention mechanism. Then, we apply token shifting and compute all $\bm{r}_{j}$ for $j\leq N-1$. Next, using $\phi^{\text{MA}}_{q}(\bm{q}^{\text{MA}}_{t})$, $\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})$, and $\bm{r}_{j}$, we calculate $\bm{o}^{\text{MA}}_{t}$ with the efficient method matching AR attention, as illustrated in Fig. 3. Finally, the ARMA output is $\bm{o}_{t}=(\bm{o}^{\text{AR}}_{t}+\bm{o}^{\text{MA}}_{t})\mathbf{W}_{o}$. A summary of MA computation methods for each attention mechanism is in Table 1.
Computational cost and model performance   The introduction of MA term adds three weight matrices $\mathbf{W}^{\text{MA}}_{\{q,k,v\}}$, increasing parameter size. To ensure fair comparison, we use weight-sharing to match the parameter sizes of ARMA and AR models. Specifically, we share $\mathbf{W}_{q}$ between the AR and MA terms and set $\mathbf{W}_{v}$ to an identity matrix, with minimal impact due to the existance of $\mathbf{W}_{o}$ and the MLP layer (see Eq. 1). This reduces ARMA’s trainable weights to $\mathbf{W}_{q},\mathbf{W}^{\text{AR}}_{k},\mathbf{W}^{\text{MA}}_{k},\mathbf{W}_{o}$, as shown in Fig. 3. While ARMA attention has the same time complexity in order of magnitude as efficient AR attention, its two-stage structure may increase computational costs on constant level. We compare models with different number of layer in the experiements section to show that ARMA’s improved performance is due to structural enhancements, but not increased complexity.
Table 1: Summary of ARMA attention for various attention mechanisms, detailing the calculation methods for AR output and MA output, where $\bm{r}_{j}=\bm{v}_{j+1}-\bm{o}^{\text{AR}}_{j}$.
@@ 2.7 Selection of $\phi(\cdot)$ and characteristics of implicit MA weights
3 Experiments
We conducted comprehensive experiments on 12 widely-used TSF datasets, including Weather, Solar, Electricity (ECL), ETTs, Traffic, and PEMS. See §A.1 for detailed description of datasets.
Baselines   We built AR Transformers using the five attention mechanisms from Table 1 and added MA terms to create ARMA attention for comparison in TSF tasks. Additionally, we included five recent SOTA baselines: FITS (Xu et al., 2024), iTransformer (Liu et al., 2024a), CATS (Lu et al., 2024), PatchTST (Nie et al., 2022), and DLinear (Zeng et al., 2023). We also used a simple channel-dependent encoder-only Transformer, modified by repeating the last input value (like NLinear) to address distribution shift. This model already surpasses older architectures like Autoformer (Wu et al., 2021) and Informer (Zhou et al., 2021), so we excluded these from our comparison.
In the main experiments, both AR and ARMA Transformers use a consistent setup: $m=3$ Transformer layers, 8 heads, and model dimension determined by a empirical method $d=16\sqrt{C}$, where $C$ is the number of series. We evaluate their performance using one-step prediction for each test datapoint, aligned with the baselines. Baseline hyperparameters are set to the reported values from their original papers. For more details on hyperparameters and implementation, see §A.2.
$L_{I}$, $L_{P}$, and fair comparison   AR/ARMA Transformers maintain stable performance across different input lengths $L_{I}$ and forecasting horizons $L_{P}$. In the main experiments, we set $L_{I}=512$ and test with $L_{P}\in\{12,24,48,96\}$. We did not select longer $L_{P}$ values for the main experiments because the token length in our AR/ARMA Transformer is determined by $L_{I}/L_{P}$. For longer horizons like $L_{P}=720$, AR/ARMA Transformers need a larger $L_{I}$ (e.g., 4096). If $L_{I}=512$ is used, it results in only one token, reducing the model to a simple MLP, which is feasible but fails to demonstrate the strengths of the AR/ARMA structure. However, increasing $L_{I}$ for baselines would hurt their performance, as $L_{I}=512$ is within the optimal range for them. To ensure fairness, we use these settings for the main experiments and later include additional experiments extending $L_{I}$ to 4096 and $L_{P}$ to 720.
We ran all models on all datasets for the four different $L_{P}$. In the main text, we report the average test set MSE for each model across different $L_{P}$ on each dataset and provide the full results in §A.3.
Main TSF results   Table 2 highlights the significant performance gains from extending AR Transformers to the ARMA structure. All ARMA attention mechanisms outperform their AR counterparts in both average test MSE and ranking, with linear and standard attention showing the best results.
Performance of gated linear attention   ARMA brought the greatest improvement to gated linear attention. In gated AR models, the decay factor helps the AR term focus on important local patterns, but it weakens the ability to capture long-term or stable cyclic patterns. By introducing the MA term, local effects are absorbed, allowing the decay factor to function properly in the AR forgetting mechanism, leading to significant performance gains.
Performance of fixed attention   Fixed attention, which lacks dynamic parameter generation, performs worse than other attention. However, its significant improvement with MA terms shows that ARMA enhances the model’s ability structurally to capture comprehensive sequence patterns.

### v2

# WAVE: Weighted Autoregressive Varing Gate for Time Series Forecasting

@@ Abstract
We propose a Weighted Autoregressive Varing gatE (WAVE) attention mechanism equipped with both Autoregressive (AR) and Moving-average (MA) components. It can adapt to various attention mechanisms, enhancing and decoupling their ability to capture long-range and local temporal patterns in time series data. In this paper, we first demonstrate that, for the time series forecasting (TSF) task, the previously overlooked decoder-only autoregressive Transformer model can achieve results comparable to the best baselines when appropriate tokenization and training methods are applied. Moreover, inspired by the ARMA model from statistics and recent advances in linear attention, we introduce the full ARMA structure into existing autoregressive attention mechanisms. By using an indirect MA weight generation method, we incorporate the MA term while maintaining the time complexity and parameter size of the underlying efficient attention models. We further explore how indirect parameter generation can produce implicit MA weights that align with the modeling requirements for local temporal impacts. Experimental results show that WAVE attention that incorporates the ARMA structure consistently improves the performance of various AR attentions on TSF tasks, achieving state-of-the-art results. The code implementation is available at the following link.
Machine Learning, ICML
@@ 1 Introduction
In recent years, autoregressive (AR) decoder-only Transformer-based models (Vaswani, 2017; Radford, 2018) have been widely used in sequence modeling tasks across fields such as NLP (Brown et al., 2020; Touvron et al., 2023), CV (Chen et al., 2020; Esser et al., 2021; Chang et al., 2022), and audio (Borsos et al., 2023). This structure is well-suited for various sequential generation and prediction tasks. However, in typical sequence modeling tasks like time series forecasting (TSF), there has been less exploration of this architecture compared to other structures. Most of the best-performing recent TSF models are encoder-only Transformers (Liu et al., 2024a; Nie et al., 2022), MLPs (Das et al., 2023; Lu et al., 2024), or even linear models (Zeng et al., 2023; Xu et al., 2024). The few relevant discussions mainly focus on using pretrained autoregressive LLMs or similar structures for few-shot and zero-shot prediction (Gruver et al., 2023; Jin et al., 2024; Das et al., 2024; Liu et al., 2024b), with little research directly evaluating their TSF performance in end-to-end training. Therefore, this paper will first briefly demonstrate that with appropriate tokenization and training methods, a basic AR Transformer is enough to achieve results comparable to the state-of-the-art (SOTA) baselines, as shown in Fig. 1.
Recently, efficient linear autoregressive attention variants have been explored and developed (Katharopoulos et al., 2020; Hua et al., 2022), reducing the time complexity of standard softmax attention from $O(N^{2})$ to $O(N)$. Researchers have found that adding a gating decay factor or a similar exponential moving average (EMA) structure to AR structure, as in gated linear attention (Ma et al., 2022; Yang et al., 2024), enhances linear attention’s ability to model local patterns and improves performance. The success of these approaches inspired us to introduce a more comprehensive full autoregressive moving-average (ARMA) structure into existing AR attention mechanisms and explore the performance of these Transformers in TSF.
@@
Figure 1: (Left: a) Overall architecture of our decoder Transformer for TSF. (Right: b) Box plots of performance rankings from 48 sub-experiments across 12 datasets. Green represents WAVE Transformers, yellow AR Transformers, and red the baselines, with triangles indicating mean rankings. AR Transformers perform comparably to baselines, while WAVE Transformers significantly outperform their AR counterparts. See Table and for more details.
We propose the Weighted Autoregressive Varing gatE (WAVE) attention mechanism equipped with ARMA structure, which integrates a moving-average (MA) term into various existing AR attention mechanisms. Our method improves the TSF performance of AR Transformers without significantly increasing computational costs, maintaining $O(N)$ time complexity and original parameter size. We design an indirect MA weight generation to obtain the MA output without explicitly computing the MA attention matrix, preserving efficiency of linear attentions. We explore specific techniques for generating implicit MA weights to ensure proper decoupling and handling of short-term effects. Extensive experiments and visualization analyses demonstrate that ARMA balances long- and short-term dependencies, significantly improving AR Transformers and achieving state-of-the-art TSF results.
@@
a) We demonstrate that, with appropriate tokenization and preprocessing methods, an AR Transformer is enough to achieve the level of existing SOTA baselines. Furthermore, the introduction of WAVE attention enables the decoder-only Transformer to outperform SOTA baselines.
b) We propose the WAVE attention mechanism, which introduces an MA term into existing AR attention without increasing time complexity or parameter size. By adding the MA term to various AR attention mechanisms, the resulting WAVE Transformers significantly improve forecasting performance compared to their AR counterparts.
@@ 2.2 Appropriate tokenization for autoregressive forecasting
To prevent error accumulation, we use an autoregressive Transformer (Fig. 1) that treats one-step prediction as the complete forecast. Inspired by PatchTST (Nie et al., 2022), we adopt a channel-independent approach, predicting each series separately and applying RevIN (Kim et al., 2022) to each. For an input series of length $L_{I}$ in $\mathbf{S}_{I}$, we apply non-overlapping patches with a patch size $L_{P}$, dividing the input into $N=\frac{L_{I}+P}{L_{P}}$ patches, where $P$ is zero-padding for divisibility. This ensures that each out-of-sample prediction token covers the entire forecasting length $L_{P}$, thereby avoiding error accumulation [^1] Please note that, consistent with previous studies on long-term time series forecasting, this paper focuses on one-step prediction for the next long time period. Specifically, we predict the next token of length $L_{P}$. Readers can view AR tokenization as PatchTST-style tokenization with an added autoregressive loss..
@@ 2.3 Preliminaries: decoder-only Transformer
$\begin{gathered}\text{Attn}(\mathbf{X})=\text{softmax}\left(\mathbf{M}\odot(\mathbf{Q}\mathbf{K}^{\top})\right)\mathbf{V}\mathbf{W}_{o},\\ \text{with}\ \ \mathbf{Q},\mathbf{K},\mathbf{V}=\mathbf{X}\mathbf{W}_{q},\mathbf{X}\mathbf{W}_{k},\mathbf{X}\mathbf{W}_{v}\\ \mathbf{X}:=\mathbf{X}+\text{Attn}(\text{LN}(\mathbf{X})),\ \ \text{then}\ \ \mathbf{X}:=\mathbf{X}+\text{MLP}(\text{LN}(\mathbf{X}))\end{gathered}$
@@ 2.4 Preliminaries: efficient linear attention mechanisms
2.5 Inspiration: decoupling the short-term impact
In sequence modeling tasks like NLP, context tokens closer to the output token typically carry higher importance. As a result, a gating mechanism with exponential decay in gated linear attention can significantly enhance the performance by assigning greater weights to nearby tokens. Additionally, NLP tasks require retrieval of long-term information. Even though the decay factor reduces the weights for long-term tokens, the AR weights’ ability to capture long-term dependencies allows for effective retrieval within moderately sized lookback windows.
However, applying exponential decay to the gated linear attention AR weights may not align with the needs of TSF, which often involves stable periodic patterns alongside short-term impacts. TSF data frequently exhibit seasonal effects that differ from the transient long-term effects in NLP. These seasonal effects are stable and persist across the temporal dimension without decaying. As shown in Figure LABEL:fig:exp_decay_strategies, in scenarios involving seasonal effects, sudden changes, and local effects, relying solely on exponential decay in gated linear attention is not the most suitable approach for modeling TSF. Decoupling local effects into short-term MA weights, allowing the AR term to focus on modeling the seasonal and long-term effects it handles best, would be a better solution for TSF.
2.6 WAVE attention mechanism
Figure 4: WAVE attention structure with the indirect MA weight generation method applied to softmax and linear attention. See Table 1 for more calculation details.
In the attention mechanisms above, the next-step prediction at time $t$ is a weighted sum of all previous values $\bm{v}_{i}\in\mathbb{R}^{1\times d}$, with weights $\mathbf{w}_{t,i}\in\mathbb{R}^{1\times d}$ derived from interactions between $\bm{q}_{t}$ and $\bm{k}_{i}$. Naturally, we can write these attention mechanisms in an AR model structure:
@@ 2.6 WAVE attention mechanism
$\bm{v}_{t+1}$ $=\bm{o}^{\text{AR}}_{t}+\bm{o}^{\text{MA}}_{t}+\bm{\epsilon}_{t}$ (2)
$=\sum_{i=1}^{t}\mathbf{w}_{t,i}\odot\bm{v}_{i}+\sum_{j=1}^{t-1}\bm{\theta}_{t-1,j}\odot\bm{\epsilon}_{j}+\bm{\epsilon}_{t},$
$\bm{r}_{t}$ $=\sum_{j=1}^{t-1}\bm{\theta}_{t-1,j}\odot\bm{\epsilon}_{j}+\bm{\epsilon}_{t}$
The structure of the MA output $\bm{o}^{\text{MA}}_{t}=\sum_{j=1}^{t-1}\bm{\theta}_{t-1,j}\odot\bm{\epsilon}_{j}$ resembles the AR term and could potentially be computed using an attention mechanism. For simplicity, we consider a single channel of the $d$-dimensional space, with other channels an be handled in parallel. We express the matrix form of the $\bm{r}_{t}$ in Eq. (2) for one channel as:
$\begin{gathered}\begin{pmatrix}r_{1}\\ r_{2}\\ \vdots\\ r_{t}\end{pmatrix}=\begin{pmatrix}0&0&\cdots&0&0\\ \theta_{1,1}&0&\cdots&0&0\\ \theta_{2,1}&\theta_{2,2}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \theta_{t-1,1}&\theta_{t-1,2}&\cdots&\theta_{t-1,t-1}&0\end{pmatrix}\begin{pmatrix}\epsilon_{1}\\ \epsilon_{2}\\ \vdots\\ \epsilon_{t}\end{pmatrix}+\begin{pmatrix}\epsilon_{1}\\ \epsilon_{2}\\ \vdots\\ \epsilon_{t}\end{pmatrix}\\ \mathbf{r}=(\mathbf{I}+\mathbf{\Theta})\bm{\epsilon},\ \ \bm{\epsilon}=(\mathbf{I}+\mathbf{\Theta})^{-1}\mathbf{r},\end{gathered}$
where $\mathbf{\Theta}$ is a strictly lower triangular matrix of MA weights for this channel. Once the attention mechanism determines $\bm{o}^{\text{AR}}_{t}$ and $\bm{\theta}_{t-1,j}$, we can calculate $\bm{r}_{j}=\bm{v}_{j+1}-\bm{o}^{\text{AR}}_{j}$ (token shifting) for all $j\leq N-1$ and determine $\epsilon_{j}$ via matrix inversion. Substituting these back into Eq. (2) yields the final WAVE attention output $\bm{o}^{\text{AR}}_{t}+\bm{o}^{\text{MA}}_{t}$ for step $t$.
@@
Table 1: Summary of WAVE attention for various attention mechanisms, detailing the calculation methods for AR output and MA output, where $\bm{r}_{j}=\bm{v}_{j+1}-\bm{o}^{\text{AR}}_{j}$.
@@
2.7 Indirect MA weight generation
We need an approach that can leverage linear attention’s efficiency to compute $\bm{o}^{\text{MA}}_{t}$ without the costly $\mathbf{\Theta}^{N\times N}$ matrix operations. Instead of separately calculating attention weights to determine $\bm{\epsilon_{j}}$ as value input and recomputing the whole MA output, we aim to use a linear RNN to collect all keys and values at once. We observe from Eq. (3) that there is already a sequential relationship between $\bm{r}_{j}$ and $\bm{\epsilon}_{j}$, and $\bm{r}_{j}$ can be computed directly once $\bm{o}^{\text{AR}}_{t}$ is determined. Therefore, we implicitly compute the MA weights of $\bm{\epsilon_{j}}$ by using $\bm{r}_{j}$ as value input for the MA component instead of $\bm{\epsilon}_{j}$. Let $\bm{\beta}_{t-1,j}$ denote the generated attention weights corresponding to $\bm{r}_{j}$ at step $t$, and let $\bm{\theta}_{t-1,j}$ here be the implicit MA weights hiddenly linked to the generated $\bm{\beta}_{t-1,j}$. Based on Eq. (3), for each channel, we establish:
$\begin{gathered}\sum_{j=1}^{t-1}\bm{\beta}_{t-1,j}\odot\bm{r}_{j}=\sum_{j=1}^{t-1}\bm{\theta}_{t-1,j}\odot\bm{\epsilon}_{j}\Leftrightarrow\mathbf{B}\mathbf{r}=\mathbf{\Theta}\mathbf{\epsilon}\\ \mathbf{B}=\mathbf{\Theta}\cdot(\mathbf{I}+\mathbf{\Theta})^{-1},\ \ \mathbf{\Theta}=\mathbf{B}\cdot(\mathbf{I}-\mathbf{B})^{-1}\end{gathered}$ (4)
$\theta_{ij}=b(1+b)^{i-j-1},\quad\text{where}\ i>j$ (5)
We tested various activation function combinations for $\phi_{q}^{\text{MA}}(\cdot)$ and $\phi_{k}^{\text{MA}}(\cdot)$ to generate $\bm{\beta}_{t-1,j}=\phi_{q}^{\text{MA}}(\bm{q}^{\text{MA}}_{t-1})\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})$ values, as shown in Fig. LABEL:fig:vis_phi_selection. We used the sigmoid function $\phi_{k}^{\text{MA}}(\bm{k}^{\text{MA}}_{j})=\sigma(\alpha\bm{k}^{\text{MA}}_{j}/\sqrt{d})$ to obtain values between 0 and 1, where $\alpha=0.05$  [^2] In the key activation, $\alpha$ controls the variance of each row in the $\mathbf{B}$ matrix, indirectly influencing the amount of long-term information (lower left) in the MA weights $\mathbf{\Theta}$. Increasing $\alpha$ would make the MA weights focus more on modeling long-term information. However, since we want the AR weights to handle the long-term component, we set $\alpha$ to a relatively small value. This explains why the rows of the $\mathbf{B}$ matrix appear smooth in the visualization. Refer to Fig. LABEL:fig:vis_phi_selection_alpha_selection for more details on $\alpha$, and see Fig. LABEL:fig:vis_phi_selection_positive for the effects of reversed positive $\phi_{q}$. and $\sqrt{d}$ are scaling factors to maintain small absolute values. Then, we selected a function $\phi_{q}^{\text{MA}}(\cdot)$ to make the product negative. We ultimately chose $\phi^{\text{MA}}_{q}(\bm{q}^{\text{MA}}_{t})=-\text{LeakyReLU}(-\bm{q}^{\text{MA}}_{t}/\sqrt{d})$ with a negative slope of 0.02. The inner negative sign maintains directional consistency (for later parameter sharing), and the outer negative sign encourages a negative output.